an 18% iodine solution is mixed with a 45% iodine solution to make 12 ounces of a 36% iodine solution. How many ounces of each were used?
Answer by lwsshak3(11628)

an 18% iodine solution is mixed with a 45% iodine solution to make 12 ounces of a 36% iodine solution. How many ounces of each were used?
**
let x= amt of 18% iodine solution to mix.
12-x= amt of 45% iodine solution to mix.
..
18%x+45%(12-x)=36%*12
.18x+5.4-.45x=4.32
.27x=1.08
x=1.08/.27
x=4
12-x=8
amt of 18% iodine solution to mix=4 ounces
amt of 45% iodine solution to mix=8 ounces